# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ 6.2.2 Results and Analysis
RTD(SEARLE)24’Arxiv CLIP(ViT-L/14) $16.53$ $17.89$ $19.77$ $20.68$
RTD(LinCIR)24’Arxiv CLIP(ViT-L/14) $17.11$ $18.11$ $20.06$ $21.01$
ISA(Sym)24’ICLR BLIP $9.67$ $10.32$ $11.26$ $11.61$
ISA(Asy)24’ICLR EfficientNet+BLIP $11.33$ $12.25$ $13.42$ $13.97$
Slerp24’ECCV CLIP(ViT-B/32) $6.35$ $7.11$ $8.12$ $8.75$
Slerp+TAT24’ECCV CLIP(ViT-B/32) $9.34$ $10.26$ $11.65$ $12.33$
Slerp+TAT24’ECCV CLIP(ViT-L/14) $18.46$ $19.41$ $21.43$ $22.41$
Slerp+TAT24’ECCV BLIP(ViT-L/16) $17.84$ $18.44$ $20.24$ $21.07$
LDRE24’SIGIR CLIP(ViT-B/32) $17.96$ $18.32$ $20.21$ $21.11$
LDRE24’SIGIR CLIP(ViT-L/14) $23.35$ $24.03$ $26.44$ $27.50$
LDRE24’SIGIR CLIP(ViT-G/14) $31.12$ $32.24$ $34.95$ $36.03$
HyCIR(Pic2Word-CLIP)24’Arxiv CLIP $14.12$ $15.02$ $16.72$ $17.56$
HyCIR(Pic2Word-BLIP)24’Arxiv BLIP $18.91$ $19.67$ $21.58$ $22.49$
@@
MCL(OPT-6.7B)21’ICML $15.14$ $16.13$ $17.88$ $18.82$
MCL(Llama2-7B)21’ICML $17.67$ $18.86$ $20.80$ $21.68$
CompoDiff(ST18M)23’TMLR CLIP(ViT-L/14) $12.55$ $13.36$ $15.83$ $16.43$
CompoDiff(ST18M)23’TMLR CLIP(ViT-G/14) $15.33$ $17.71$ $19.45$ $21.01$
TABLE VII: Performance Comparison of Different Methods on MIT-States and Birds-to-Words Datasets
@@
TIRG18’CVPR ResNet18 LSTM $12.20$ $31.90$ $43.10$ $29.07$
LBF(small)20’CVPR Faster-RCNN $14.72$ $35.3$ $46.56$ $32.19$
LBF(big)20’CVPR Faster-RCNN $14.72$ $35.3$ $46.56$ $32.19$
MAAF20’Arxiv ResNet50 LSTM $12.70$ $32.60$ $44.80$ $30.03$ $34.75$ $66.29$
TRACE21’AAAI ResNet50 GRU $19.56$ $45.24$
GSCMR21’TIP Faster-RCNN BiGRU $17.28$ $36.45$ $47.04$ $33.59$
ComposeAE21’WACV ResNet18 BERT $13.90$ $35.30$ $47.90$ $32.37$
MAN21’ICASSP ResNet18 BERT $13.90$ $35.30$ $47.90$ $31.53$
MAN21’ICASSP MobileNet BERT $15.60$ $36.70$ $47.70$ $33.33$
RTIC21’Arxiv ResNet50 LSTM $37.40$ $66.97$
HFCA21’ACMMM ResNet50 LSTM $14.30$ $35.36$ $47.12$ $32.26$
HCL21’MMAsia ResNet18 LSTM $15.22$ $35.95$ $46.71$ $32.63$
CRR22’ACMMM ResNet101 GRU $17.71$ $37.16$ $47.83$ $34.23$
TriArea22’Sci.Rep. ResNet18+TF LSTM $13.20$ $33.30$ $44.30$ $30.27$
SAC22’WACV ResNet50 GRU $19.56$ $45.24$
LIMN24’TPAMI ResNet50 LSTM $40.34$ $66.18$
@@ 6.3 Composed Video Retrieval
Composed Video Retrieval (CoVR) aims to retrieve specific videos from a large database by integrating both visual queries and modification texts, thus allowing for more sophisticated video searches. This approach addresses the limitations of traditional content-based video retrieval, which relies solely on visual features and often fails to fully capture user intent or context. The main goal of CoVR is to enhance the accuracy and relevance of video search results by leveraging multi-modal information. By combining a reference video with a textual description that specifies desired modifications or conditions, users can refine their search criteria to find videos that closely match their requirements. For instance, given a query image of a fountain in Barcelona, appending the text ”during show at night” would enable users to discover videos featuring that fountain during a nighttime show, offering a more precise and personalized search experience.
CoVR has significant application value across various domains, including e-commerce, fashion, internet video search, live event discovery, and sports video retrieval. In the e-commerce and fashion sectors, it helps customers quickly locate specific product videos showcasing particular attributes or usage scenarios, thereby enhancing the shopping experience. For online video platforms, this technology supports more effective content recommendation and management systems by accurately identifying and recommending videos that align with user interests and preferences.
@@ 6.3.1 Benchmark Datasets
CIRR [42] and FashionIQ [82] are primarily focused on composed image retrieval but also serve as benchmarks for evaluating the zero-shot performance of CoVR models. These datasets provide valuable insights into how well CoVR methods can generalize to different types of data and tasks.
@@
ICQ [93] is a pioneering benchmark specifically designed for the task of localizing events in videos using multimodal queries (MQs). This dataset is created to bridge the gap between traditional natural language queries (NLQs) and the more versatile MQs, which integrate images to represent semantic queries more flexibly. ICQ includes an evaluation dataset called ICQ-Highlight, featuring synthetic reference images and human-curated queries, offering a robust testbed for this novel task. The construction of ICQ involves utilizing a combination of reference images and refinement texts to form each query, enabling a more adaptable and nuanced application scenario. The dataset evaluates model performance across four distinct reference image styles to ensure comprehensive coverage of diverse scenarios. To accommodate existing video localization models for this new task, ICQ introduces three Multimodal Query Adaptation (MQA) methods and a Surrogate Fine-tuning on pseudo-MQs strategy, collectively known as SUIT. These approaches facilitate the integration of MQs into conventional NLQ-based backbone models.
@@ 6.4 Composed Remote Sensing Image Retrieval
Traditional remote sensing image retrieval (CRSIR) typically relies on either image content or text descriptions to find matching remote sensing images. However, this approach often falls short when precise identification of geographic features (e.g., shapes, colors, densities, targets) or specific time points is required. By integrating both image and text modalities, CRSIR allows users to submit both an example image and a textual description as query inputs. This significantly enhances the query’s descriptive power, enabling the system to interpret not only the visual content of the image but also the accompanying text, thereby refining the relevance of the search results. This method provides more accurate responses to users’ customized search needs and is applicable across various domains, including environmental monitoring, urban planning, and disaster response.
CRSIR has also proven to be highly beneficial in practical applications, offering both business and social advantages. It improves the efficiency and accuracy of image retrieval while enhancing the user experience by enabling more targeted decision-making through customized products and services. For geographic researchers, CRSIR accelerates research progress, optimizes resource utilization, and fosters interdisciplinary collaboration. In fields such as environmental monitoring, agricultural management, urban planning, and disaster response, CRSIR delivers exceptional data processing capabilities and substantial support for related research. In conclusion, CRSIR technology represents a significant advancement in remote sensing data applications, with both important academic implications and vast market potential.
@@ 6.4.1 Benchmark Datasets
PATTERNCOM [94] is a large-scale, high-resolution remote sensing image retrieval dataset consisting of 38 classes, with each class containing 800 images of 256×256 pixels. In PATTERNCOM, specific categories are selected to be described in the query image, and a corresponding query text defines attributes related to each class. For instance, the query image for the ”swimming pool” category is paired with text queries that describe attributes such as shape, with options like ”rectangular,” ”oval,” and ”kidney-shaped.” The dataset includes six attributes, each comprising up to four different classes. Each attribute can be associated with two to five values per class. The dataset contains a total of over 21,000 queries, with the number of positive queries per class ranging from 2 to 1345.

### v2

@@ 6.2.2 Results and Analysis
RTD(SEARLE)24’Arxiv [83] CLIP(ViT-L/14) $16.53$ $17.89$ $19.77$ $20.68$
RTD(LinCIR)24’Arxiv [83] CLIP(ViT-L/14) $17.11$ $18.11$ $20.06$ $21.01$
ISA(Sym)24’ICLR [80] BLIP $9.67$ $10.32$ $11.26$ $11.61$
ISA(Asy)24’ICLR [80] EfficientNet+BLIP $11.33$ $12.25$ $13.42$ $13.97$
Slerp24’ECCV [88] CLIP(ViT-B/32) $6.35$ $7.11$ $8.12$ $8.75$
Slerp+TAT24’ECCV [88] CLIP(ViT-B/32) $9.34$ $10.26$ $11.65$ $12.33$
Slerp+TAT24’ECCV [88] CLIP(ViT-L/14) $18.46$ $19.41$ $21.43$ $22.41$
Slerp+TAT24’ECCV [88] BLIP(ViT-L/16) $17.84$ $18.44$ $20.24$ $21.07$
LDRE24’SIGIR [76] CLIP(ViT-B/32) $17.96$ $18.32$ $20.21$ $21.11$
LDRE24’SIGIR [76] CLIP(ViT-L/14) $23.35$ $24.03$ $26.44$ $27.50$
LDRE24’SIGIR [76] CLIP(ViT-G/14) $31.12$ $32.24$ $34.95$ $36.03$
HyCIR(Pic2Word)24’Arxiv [94] CLIP $14.12$ $15.02$ $16.72$ $17.56$
HyCIR(Pic2Word)24’Arxiv [94] BLIP $18.91$ $19.67$ $21.58$ $22.49$
MoTaDual24’Arxiv [86] CLIP(ViT-L/14) $20.42$ $21.62$
InstructCIR124’Arxiv [241] CLIP(ViT-L/14) $22.32$ $23.80$ $26.25$ $27.32$
DeG25’Arxiv [140] CLIP(ViT-L/14) $13.7$ $14.9$ $16.8$ $17.7$
InstructCIR225’Arxiv [143] CLIP(ViT-L/14) $10.98$ $12.94$ $13.84$ $15.62$
TSCIR25’Arxiv [139] CLIP(ViT-L/14) $14.79$ $15.15$ $16.92$ $19.00$
SEIZE24’ACMMM [137] CLIP(ViT-B/32) $19.04$ $19.64$ $21.55$ $22.49$
SEIZE24’ACMMM [137] CLIP(ViT-L/14) $24.98$ $25.82$ $28.24$ $29.35$
PrediCIR25’CVPR [138] CLIP(ViT-L/14) $15.7$ $17.1$ $18.6$ $19.3$
CoLLM25’CVPR [142] CLIP(ViT-L/14) $20.3$ $20.8$ $23.4$
@@
MCL(OPT-6.7B)21’ICML [85] $15.14$ $16.13$ $17.88$ $18.82$
MCL(Llama2-7B)21’ICML [85] $17.67$ $18.86$ $20.80$ $21.68$
CompoDiff23’TMLR [98] CLIP(ViT-L/14) $12.55$ $13.36$ $15.83$ $16.43$
CompoDiff23’TMLR [98] CLIP(ViT-G/14) $15.33$ $17.71$ $19.45$ $21.01$
TSCIR25’Arxiv [139] CLIP(ViT-L/14) $18.37$ $19.55$ $21.64$ $22.71$
MRA-CIR25’Arxiv [150] BLIP2(ViT-L/14) $27.14$ $28.85$ $31.54$ $32.63$
CAT-LLM25’CVPR [148] CLIP(ViT-L/14) $15.00$ $15.73$ $17.51$ $18.45$
CAT-LLM25’CVPR [148] CLIP(ViT-B/16) $13.95$ $14.47$ $16.00$ $16.74$
IP-CIR(LDRE)25’CVPR [149] CLIP(ViT-L/14) $26.43$ $27.41$ $29.87$ $31.07$
IP-CIR(LDRE)25’CVPR [149] CLIP(ViT-G/14) $32.75$ $34.26$ $36.86$ $38.03$
CIG-XL(SEARLE)25’CVPR [148] CLIP(ViT-B/32) $10.3$ $10.79$ $12.12$ $12.76$
CIG-XL(LinCIR)25’CVPR [148] CLIP(ViT-L/14) $12.97$ $13.64$ $15.14$ $16.01$
ConText-CIR25’CVPR [148] CLIP(ViT-L/14) $30.05$ $30.53$ $34.79$ $34.72$
TABLE VII: Performance Comparison of Different Methods on MIT-States and Birds-to-Words Datasets.
@@
TIRG18’CVPR [20] ResNet18 LSTM $12.20$ $31.90$ $43.10$ $29.07$
LBF(small)20’CVPR [40] Faster-RCNN $14.72$ $35.3$ $46.56$ $32.19$
LBF(big)20’CVPR [40] Faster-RCNN $14.72$ $35.3$ $46.56$ $32.19$
MAAF20’Arxiv [49] ResNet50 LSTM $12.70$ $32.60$ $44.80$ $30.03$ $34.75$ $66.29$
TRACE21’AAAI [50] ResNet50 GRU $19.56$ $45.24$
GSCMR21’TIP[62] Faster-RCNN BiGRU $17.28$ $36.45$ $47.04$ $33.59$
ComposeAE21’WACV [38] ResNet18 BERT $13.90$ $35.30$ $47.90$ $32.37$
MAN21’ICASSP [136] ResNet18 BERT $13.90$ $35.30$ $47.90$ $31.53$
MAN21’ICASSP [136] MobileNet BERT $15.60$ $36.70$ $47.70$ $33.33$
RTIC21’Arxiv [58] ResNet50 LSTM $37.40$ $66.97$
HFCA21’ACMMM [42] ResNet50 LSTM $14.30$ $35.36$ $47.12$ $32.26$
HCL21’MMAsia [52] ResNet18 LSTM $15.22$ $35.95$ $46.71$ $32.63$
CRR22’ACMMM [59] ResNet101 GRU $17.71$ $37.16$ $47.83$ $34.23$
TriArea22’Sci.Rep. [44] ResNet18+TF LSTM $13.20$ $33.30$ $44.30$ $30.27$
SAC22’WACV [45] ResNet50 GRU $19.56$ $45.24$
LIMN24’TPAMI [57] ResNet50 LSTM $40.34$ $66.18$
@@ 6.3 Composed Video Retrieval
Composed Video Retrieval (CoVR) enables the retrieval of specific videos from large databases by integrating visual queries with textual modification instructions, allowing for more precise and context-aware searches. This approach overcomes the limitations of traditional content-based video retrieval systems, which rely solely on visual features and often fail to capture user intent or nuanced context. The primary objective of CoVR is to improve search accuracy by leveraging multi-modal inputs. CoVR holds strong application potential across multiple domains, including online video platforms, live event discovery, and sports video retrieval. On video platforms, it supports advanced content recommendation and management systems by identifying and suggesting videos that better align with user preferences and interests.
TABLE VIII: Performance Comparison on WebVid-CoVR-Test Dataset.
Method Backbone WebVid-CoVR-Test
R@1 R@5 R@10 R@50
CoVR(Avg)24 [29] CLIP 44.37 69.13 77.62 93.00
CoVR(Avg)24 [29] BLIP 45.46 70.46 79.54 93.27
CoVR-2(Avg)24 [108] BLIP-2 45.66 71.71 81.30 94.80
CoVR(ft-CA)24 [29] BLIP 53.13 79.93 86.85 97.69
CoVR-2(ft-CA)24 [108] BLIP 55.95 81.22 89.05 98.08
CoVR-2(ft-CA)24 [108] BLIP-2 59.82 83.84 91.28 98.24
ECDE(ft-CA)24 [108] BLIP 60.12 84.32 91.27 98.72
TABLE IX: Performance Comparison on EgoCVR Dataset.
Method Backbone EgoCVR
Global Local
R@1 R@5 R@10 R@1 R@5 R@10
BLIP-CoVR24 [29] BLIP 5.4 15.2 24.3 33.1 49.5 62.9
BLIP-CoVR-ECDE24 [84] BLIP 6.0 16.3 24.8 33.4 49.3 63.0
CIReVL24 [74] CLIP 2.0 6.8 10.6 33.6 49.7 61.4
TFR-CVR24 [30] BLIP 14.1 39.5 54.4 44.2 61.0 73.2
TABLE X: Performance Comparison on Airplane, Tennis, and WHIRT.
Method Datasets
Airplane Tennis WHIRT
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
TIRG19 [20] 15.05 37.28 52.20 5.68 22.16 35.04 0.83 4.18 4.56
ComposeAE21 [38] 16.18 45.34 61.40 5.90 23.58 37.55 1.62 5.17 9.58
Cosmo21 [37] 20.91 49.31 62.66 11.68 39.63 55.02 1.71 7.23 13.15
CLIP4Cir22 [33] 20.53 51.19 63.85 17.58 46.51 61.68 2.10 9.86 17.62
AACL23 [246] 14.29 45.53 61.08 7.42 28.71 45.41 2.28 7.93 14.64
UncerRe22 [101] 19.14 51.13 66.25 13.97 43.78 60.59 1.84 9.16 17.66
SHF24 [110] 62.15 96.22 98.43 43.23 84.93 93.89 5.08 21.60 36.33
TABLE XI: Performance Comparison on PATTERNCOM Dataset.
Method PATTERNCOM
Color Context Density Existence Quantity Shape Avg
WEICOM24 [109]
(CLIP) 46.74 20.97 22.07 12.07 20.96 26.22 24.83
WEICOM24 [109]
(RemoteCLIP) 41.04 31.59 41.56 14.79 20.79 31.24 30.19
TABLE XII: Performance Comparison on ITCPR Dataset.
Method (Pre-trained on SynCPR) ITCPR
R@1 R@5 R@10 mAP
CaLa24 [60] 39.33 60.85 68.66 49.29
SPRC24 [63] 42.27 61.81 69.35 51.62
FAFA25 [111] 46.54 66.21 73.12 55.60
TABLE XIII: Performance Comparison on Sketch-based Image Dataset.
Method ShoeV2 ChairV2 Sketchy FS-COCO SketchyCOCO
R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10
Combiner22 [33] 24.7 40.2 35.7 39.9 15.7 33.7 11.6 22.1 15.9 32.2
TASK-former22 [247] 27.7 44.1 40.7 45.2 17.8 35.2 12.7 24.2 19.4 34.7
SceneTrilogy23 [248] 29.1 46.2 43.4 46.8 19.7 37.2 14.5 28.3 20.4 40.2
Pic2Word23 [12] 34.7 58.4 55.7 62.1 22.5 48.7 16.7 32.6 24.4 46.0
SEARLE23 [23] 38.4 64.8 60.8 66.4 25.3 54.2 17.7 35.9 26.0 50.4
SCIR24 [112] 47.3 79.1 73.5 81.4 30.6 64.2 22.7 43.5 33.4 61.1
TABLE XIV: Performance Comparison on Multiturn-Fashion-IQ test set.
Method Multiturn-Fashion-IQ test set
Dress Shirt Toptee
R@5 R@8 MRR R@5 R@8 MRR R@5 R@8 MRR
TIRG19 [20] 12.5 14.2 11.6 13.8 16.7 12.9 12.0 15.6 10.9
RTIC20 [249] 11.8 18.8 10.2 14.0 20.6 12.2 13.2 18.9 11.7
ComposeAE21 [38] 18.5 26.4 14.4 19.8 25.2 14.5 19.2 26.6 14.8
CCNet20 [250] 12.7 17.2 10.5 15.2 18.5 13.3 13.6 16.2 12.1
AUS21 [26] 13.4 15.3 10.5 14.7 16.6 11.3 12.4 13.3 10.6
Dialog Manager18 [28] 12.7 16.7 10.8 13.9 17.7 11.6 11.6 15.8 10.3
IRR23 [118] 26.8 31.2 20.6 25.8 30.4 19.8 27.1 31.7 20.9
CAFA21 [115] 30.3 33.4 26.5 29.8 33.5 25.6 30.5 34.1 27.4
FashionNTM23 [117] 48.3 52.8 45.1 49.8 43.8 48.8
@@ 6.3.1 Benchmark Datasets
CIRR [21] and FashionIQ [26] are primarily focused on composed image retrieval but also serve as benchmarks for evaluating the zero-shot performance of CoVR models. These datasets provide insights into how well CoVR methods can generalize to different types of data and tasks.
@@
ICQ [31] is a pioneering benchmark specifically designed for the task of localizing events in videos using multimodal queries (MQs). ICQ includes an evaluation dataset called ICQ-Highlight, featuring synthetic reference images and human-curated queries, offering a robust testbed for this novel task. The dataset evaluates model performance across four distinct reference image styles to ensure comprehensive coverage of diverse scenarios.
@@ 6.3.2 Results and Analysis
We report experimental results for composed video retrieval on two datasets: WebVid-CoVR-Test (Table VIII) and EgoCVR (Table IX). On the WebVid-CoVR-Test dataset, models fine-tuned with context-aware (CA) strategies show strong performance, particularly ECDE and CoVR-2(ft-CA), with Recall@10 exceeding 90% and Recall@1 reaching up to 60%. These results demonstrate the effectiveness of context modeling and pretraining strategies in open-domain video retrieval.
In contrast, EgoCVR presents a more challenging setting that emphasizes understanding fine-grained temporal variations in egocentric videos. The task includes two subtasks: Global (retrieving the correct video from a large gallery) and Local (identifying the relevant segment within a known video). Across all methods, performance on the Local task is consistently higher than on the Global task, highlighting the inherent difficulty of coarse-level retrieval across semantically similar long videos. In this setting, TFR-CVR achieves a notable gain in the Global subtask, improving Recall@1 from 5.4% to 14.1%, while the improvement in the Local task is relatively modest (from 33.1% to 44.2%). This asymmetric gain indicates that the model excels in video-level semantic discrimination. Nonetheless, even the best-performing models still fall short in precise temporal localization, suggesting that current approaches are limited in capturing fine-grained temporal dependencies.
In summary, although models in the CoVR line of work demonstrate strong performance on object-centric, open-domain benchmarks such as WebVid, substantial improvements are still needed for temporally grounded and semantically subtle scenarios like EgoCVR, particularly in fine-grained time reasoning and cross-frame semantic alignment.
@@ 6.4 Composed Remote Sensing Image Retrieval
Composed Remote Sensing Image Retrieval (CRSIR) enables users to perform more precise and expressive searches by combining both visual and textual inputs. Instead of relying on a single modality, users can submit a reference image together with a textual description that specifies desired geographic features, environmental conditions, or temporal information. This multi-modal approach enhances the system’s ability to interpret complex queries, leading to more accurate and context-aware retrieval results.
CRSIR is particularly valuable in applications that require high levels of specificity and customization. In environmental monitoring, for instance, it allows users to search for images of a particular landscape with added conditions such as vegetation type, season, or land use changes. In urban planning, the method supports detailed analysis of infrastructure development across time and space. In conclusion, CRSIR technology represents a significant advancement in remote sensing data applications, with both important academic implications and vast market potential.
@@ 6.4.1 Benchmark Datasets
PATTERNCOM [109] is a large-scale, high-resolution remote sensing image retrieval dataset consisting of 38 classes, with each class containing 800 images of 256×256 pixels. In PATTERNCOM, specific categories are selected to be described in the query image, and a corresponding query text defines attributes related to each class. For instance, the query image for the ”swimming pool” category is paired with text queries that describe attributes such as shape, with options such as ”rectangular”, ”oval”, and ”kidney-shaped.” The dataset includes six attributes, each comprising up to four different classes. Each attribute can be associated with two to five values per class. The dataset contains a total of over 21,000 queries, with the number of positive queries per class ranging from 2 to 1345.
@@ 6.4.2 Results and Analysis
We report composed remote sensing image retrieval results on two task subsets: the Airplane, Tennis, and WHIRT datasets (Table X), and the PATTERNCOM dataset (Table XI). The results in Table X clearly demonstrate the substantial impact of scene complexity on retrieval performance. On the relatively simple and well-defined Airplane and Tennis datasets, early generic methods such as TIRG and CLIP4Cir achieve only around 15%–20% in Recall@1. In contrast, SHF, a method specifically designed for remote sensing imagery, shows a significant performance gain, achieving Recall@1 scores of 62.15% and 43.23% on Airplane and Tennis, respectively. This improvement suggests that conventional composition strategies are inadequate for remote sensing tasks, where architectures like SHF, capable of explicitly modeling spatial structure and hierarchical semantics, are essential. Conversely, all methods exhibit a marked performance drop on the more diverse and complex WHIRT dataset. Even SHF achieves only 5.08% in Recall@1, underscoring the considerable challenge of generalizing to more heterogeneous remote sensing scenes.
Further experimental results indicate that generalization in remote sensing depends not only on model architecture but also on the incorporation of domain-specific knowledge during pretraining. On the PATTERNCOM dataset (Table XI), the WEICOM model improves its average recall from 24.83% to 30.19% when the general-purpose CLIP encoder is replaced with RemoteCLIP, a backbone pretrained on remote sensing data. These findings highlight the importance of domain-adapted pretraining for capturing unique visual characteristics and enhancing vision-language alignment.